OpenStudy (xenesis98):

MEDALS!Which expression is equivalent to (1 – sin^2 x)(sec x)? sec^2 x tan x tan^2 x cos x cos^2 x

OpenStudy (salazarblack):

are you familiar with wolfram alpha?

OpenStudy (xenesis98):

nope

OpenStudy (xenesis98):

its a random online thingy they dont teach

OpenStudy (luigi0210):

You can change \(\large (1-sin^2 x)\) into \(\large cos^2 x \) and change \(\large sec x \) into \(\large \frac{1}{cosx} \). That should help .

OpenStudy (salazarblack):

they don't but you can see how to simplify things

OpenStudy (error1603):

Use sin² x + cos² x = 1 → so replace 1 - sin² x with cos² x and sec x = 1 / cos so (1 - sin² x)(sec x) = (cos² x) * (1 / cos x) = cos x →→ so choice is ?

OpenStudy (xenesis98):

d`?

OpenStudy (xenesis98):

is it d?

OpenStudy (salazarblack):

no

OpenStudy (xenesis98):

idk then e?

OpenStudy (salazarblack):

yes it's d ..sorry i thought d was the last one

OpenStudy (xenesis98):

oh thanks